OpenStudy (anonymous):

Use the technique of completing the square to transform the quadratic equation into the form (x + c)2 = a. 6x2 + 36x + 18 = 0

OpenStudy (btaylor):

first, take out a gcf of 6 x2+6x+3=0 move the constant to the other side: x2+6x=-3 make a square (6=2 x 3, so it is 3^2) x2+6x+9=-3+9 (x+3)^2=6 get it?

OpenStudy (anonymous):

This is basically college work, and I'm still in High School, we've never talked about this, so I don't understand it.